OpenStudy (anonymous):

The U.S.A. has the world's only "large deck" aircraft carriers which can hold up to 72 aircrafts. The Enterprise class carrier is longest in length while the Nimitz class carrier is the second longest. The total length of these two carriers is 2193 feet while the difference of their lengths is only 9 feet. a.) Find the length of each class carrier b.) If a football field has a length of 100 yards, determine the length of the Enterprise class carrier in terms of number of football fields.

OpenStudy (wwe123):

Let x = the length of the shorter ship. x+9 = length of the longer ship <== enterprise class 2x + 9 = 2193 2x = 2184 x = 1092 x + 9 = 1101 <=== length of enterprise calss ... but in feet 1101 / 3 / 100 = 3.67 foot ball fields

OpenStudy (wwe123):

If the Enterprise class has length E and the Nimitz class has length N, then E + N = 2193 E - N = 9 and you can solve the simultaneous equations. Skyhawk's answer is wrong, by the way.

OpenStudy (anonymous):

@wwe123 Thank you for helping. =)

OpenStudy (anonymous):

@wwe123 Sorry, the webpage froze up. I was trying to say thank you for helping, but your answer is wrong.

OpenStudy (anonymous):

Well, not wrong, but it wasn't done correctly.

OpenStudy (anonymous):

This question requires a system of solutions. It should be {x=y+9, x+y=2193}. The only thing is, kids come on this site for genuine help. All you did was copy and paste from Yahoo Answers, which is a highly unreliable, sloppy source.